OpenStudy (anonymous):

brandon has made an 87, 95,86, and a 88 on his caculus test. what grade will be made on his 5th test in order to have a 90 average in his caculus class? show work

OpenStudy (inkyvoyd):

(87+95+86+88+x)/5=90

OpenStudy (dcolley):

94: Because 87+95+86+88=356+94=450/5=90

OpenStudy (inkyvoyd):

Does that equation make sense to you?

OpenStudy (inkyvoyd):

@dcolley , you could've helped him out, instead of directly giving him the answer :/

OpenStudy (dcolley):

I gave an explanation

OpenStudy (inkyvoyd):

Well, you gave an explanation for checking work, but not really finding the answer.

OpenStudy (anonymous):

The average of anything is equal to the sum of the values divided by the number of values. For example, if I got a 87,62, and 91 on three of my physics tests, the average would be equal to the sum of 87+62+91 divided by number of tests (3). So, now to your question. We have four of the five values (four of the five test scores), and we want to get an average of 90. This means that the sum of all five test scores divided by five must equal 90. So to find the grade he needs on the 5th test to have a 90 average. This is represented by the following equation. \[\frac{87+95+86+88+x}{5}=90\] From here we use basic algebra to solve for x, the fifth test score.

OpenStudy (inkyvoyd):

I got a sniper and a Texer. So going to lose this one xD

OpenStudy (dcolley):

you had already gave one for solving

OpenStudy (anonymous):

@inkyvoyd the OpenStudy metagame is quite awesome :D

OpenStudy (inkyvoyd):

Well, it's ok to have two explanations, but directly giving the answer is usually looked down upon.

OpenStudy (inkyvoyd):

@Argonx16 IKR?

OpenStudy (anonymous):

so where do you get 94 from??

OpenStudy (inkyvoyd):

(87+95+86+88+x)/5=90

OpenStudy (inkyvoyd):

@ralph123 , does that equation make sense to you?

OpenStudy (anonymous):

@ralph123 you get 94 by solving \[\frac{85+95+86+88+x}{5}=90\] Do you need a step by step guide about how to go about solving it?

OpenStudy (anonymous):

yes please! :)

OpenStudy (anonymous):

So we start out with. \[\frac{87+95+86+88+x}{5}=90\] To remove the 5 on the bottom, we multiply both sides by 5 to get. \[87+95+86+88+x=450\] From here, we subtract 87 from both sides to get \[95+86+88+x=363\] Then we subtract 95 from both sides to get \[86+88+x=268\] And then we subtract 88 and 86 from both sides to end up with \[x=94\] Where x is the 5th test score he needs.

OpenStudy (anonymous):

thanks so much man!:

OpenStudy (anonymous):

No problem :)